# The Mixed Integer Trust Region Problem

### v1

@@ 4.2 Proof of Proposition 1
In the remainder of the proof, we consider the case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$. In particular, we have $h\neq 0$. Using the algorithm in Lemma 8 (with $a:=1/\lVert h\rVert_{2}^{2}$), we find a positive $l\in\mathbb{Q}$ such that $l\leq 1/\lVert h\rVert_{2}\leq(1+\epsilon)l$. We obtain $l\lVert h\rVert_{2}\leq 1$ and $l\lVert h\rVert_{2}\geq 1/(1+\epsilon)\geq 1-\epsilon$, thus $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $\mathcal{B}_{s}:=\mathcal{B}((1-\epsilon/2)lh,\epsilon/2)$. From Lemma 9, we either find a vector $s\in\mathcal{B}_{s}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{s})\leq p2^{p(p-1)/4}$. Let $\mathcal{B}_{t}:=\mathcal{B}(-(1-\epsilon/2)lh,\epsilon/2)$. From Lemma 9, we either find a vector $t\in\mathcal{B}_{t}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{t})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $s$ and $t$. Note that $\mathcal{B}_{s}$ and $\mathcal{B}_{t}$ are contained in $\mathcal{B}(0,1)$, thus $s,t$ are both feasible to Equation MITR. We also have $\mathcal{B}_{s}\subseteq\mathcal{B}(lh,\epsilon)$ and $\mathcal{B}_{t}\subseteq\mathcal{B}(-lh,\epsilon)$, thus we can use Lemma 6 and obtain
$f(s)-f(t)\geq 2\left(1-4\epsilon-\epsilon^{2}\right)\lVert h\rVert_{2}\geq\left(1-9\epsilon/2-5\epsilon^{2}/4\right)\lVert h\rVert_{2},$
where the last inequality holds for $\epsilon\in(0,1/4]$. Next, consider the subcase where Lemma 9 did not find both vectors $s$ and $t$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, where $\mathcal{B}\in\{\mathcal{B}_{s},\mathcal{B}_{t}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{2}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{2}{\epsilon}p2^{p(p-1)/4}.$
@@
5 Theorem of the alternative for Equation E-MIQP
We are now ready to get back to Equation E-MIQP. The main goal of this section is to present a theorem of the alternative for this problem: an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. To state this result, and to work with ellipsoids, it will be useful to introduce the following notation. An ellipsoid is a set of the form
@@ 5 Theorem of the alternative for Equation E-MIQP
Proposition 2 (Theorem of the alternative for Equation E-MIQP).
@@
$\left\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\right\}\in\left[\rho,\rho+28p2^{p(p-1)/4}/\epsilon\right].$
@@
5.1 Ingredients for theorem of the alternative
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $b^{1},\dots,b^{n}$ be linearly independent vectors in $\mathbb{R}^{n}$ and let $p\in\{0,\dots,n\}$. Let $B\in\mathbb{R}^{n\times n}$ with columns $b^{1},\dots,b^{n}$, let $c\in\mathbb{R}^{n}$, and let $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the map defined by $\tau(x):=B(x-c)$. Let ${d^{\prime}}\in\mathbb{R}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$. Define $d:=B^{\mathsf{T}}{d^{\prime}}$. Then, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
@@
For $j=1,\dots,n$, we have $d_{j}={b^{j}}^{\mathsf{T}}d^{\prime}$. Thus, ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer implies $d_{1},\cdots,d_{p}$ integer, and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ implies $d_{p+1}=\cdots=d_{n}=0$. We now show that for $\mathcal{S}\subseteq\mathbb{R}^{n}$ we have $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
@@ 5.2 Proof of Proposition 2
We now have all ingredients that we need to prove Proposition 2. Before giving the proof, we give an overview of the algorithm. First, we construct an affine transformation $\tau$ that maps $\mathcal{E}(c,Q)$ into an ellipsoid that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,1+\epsilon)$. We denote by Equation ${}^{\prime}$ the optimization problem obtained by applying the same affine transformation to Equation E-MIQP. Next, we consider Equation ${}^{\prime}$-w obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon)$ and by dropping the mixed integer lattice constraint. We then find an almost optimal solution $y^{\dagger}$ to Equation ${}^{\prime}$-w and use Lemma 9 to search for a close by vector $y^{\diamond}$ that is feasible to Equation ${}^{\prime}$. Next, we consider Equation E-MIQP’-s obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP’-s that are far in objective value. If we manage to find all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation ${}^{\prime}$. Otherwise, we show that $\mathcal{E}(c,Q)$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 2, which uses Lemmas 10, 11, 9, 1 and 12.
@@
Let $\epsilon^{\prime}:=\epsilon/13\in(0,1/13]$. Using the algorithm in Lemma 10 (with $\delta:=\epsilon^{\prime}$), we find a map $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\tau(x):=B(x-c)$, with $B\in\mathbb{Q}^{n\times n}$ invertible, such that
@@
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{E}(c,Q))\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})-\{Bc\},\end{split}$ (${}^{\prime}$)
where $\operatorname{\gamma}:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, and $b^{1},\dots,b^{n}$ are the columns of $B$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider the optimization problem obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon^{\prime})$ and by dropping the mixed integer lattice constraint:
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq(1+\epsilon^{\prime})^{2}.\end{split}$ (${}^{\prime}$-w)
Let $y^{\circ}$ be an optimal solution of Equation ${}^{\prime}$-w, which we do not need to compute. Using the algorithm in Lemma 11 (with $\epsilon:=\delta:=\epsilon^{\prime}$), we find a feasible solution $y^{\dagger}$ to Equation ${}^{\prime}$-w such that
@@
Let $\mathcal{B}_{\diamond}:=\mathcal{B}((1-\epsilon^{\prime}/2)y^{\dagger}/(1+\epsilon^{\prime}),\epsilon^{\prime}/2)$. From Lemma 9, we either find a vector $y^{\diamond}\in\mathcal{B}_{\diamond}\cap(\Pi_{p}(b^{1},\dots,b^{n})-\{Bc\})$, or a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}_{\diamond})\leq p2^{p(p-1)/4}$.
Consider now the mixed integer trust region problem obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})-\{Bc\}.\end{split}$ (E-MIQP’-s)
Using the algorithm in Proposition 1, we either find two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP’-s such that
$f(y^{1})-f(y^{2})\geq\left(1-9\epsilon^{\prime}/2-5{\epsilon^{\prime}}^{2}/4\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\},$ (4)
or we find a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq 2p2^{p(p-1)/4}/\epsilon^{\prime}$.
In the remainder of the proof, we consider separately two cases. First, consider the case where Lemmas 9 and 1 successfully found all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$. Note that, since $y^{\dagger}\in\mathcal{B}(0,1+\epsilon^{\prime})$, we have $y^{\dagger}/(1+\epsilon^{\prime})\in\mathcal{B}(0,1)$, thus $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(0,1)$ and $y^{\diamond}$ is feasible to Equation E-MIQP’-s. Hence, the three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$ are feasible to Equation E-MIQP’-s and to Equation ${}^{\prime}$. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation ${}^{\prime}$. From (4), we directly obtain the following lower bound on $f_{\sup}-f_{\inf}$:
$\begin{split}f_{\sup}-f_{\inf}&\geq f(y^{1})-f(y^{2})\\ &\geq\left(1-9\epsilon^{\prime}/2-5{\epsilon^{\prime}}^{2}/4\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (5)
Next, we obtain an upper bound on $f_{y}(y^{\diamond})-f_{\inf}$. First, let $g\in\mathbb{R}^{n}$ so that $y^{\diamond}=y^{\dagger}+g$. It is simple to check that $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(y^{\dagger},2\epsilon^{\prime})$, which implies $\lVert g\rVert_{2}\leq 2\epsilon^{\prime}$. Then,
@@
$\leq 4\epsilon^{\prime}\lVert H^{\prime}\rVert_{2}+4{\epsilon^{\prime}}^{2}\lVert H^{\prime}\rVert_{2}+2\epsilon^{\prime}\lVert{h^{\prime}}\rVert_{2}$
$\leq\left(6\epsilon^{\prime}+4{\epsilon^{\prime}}^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$
@@
$\begin{split}f_{y}(y^{\diamond})-f_{\inf}&\leq f_{y}(y^{\diamond})-f_{y}(y^{\circ})\\ &\leq\lvert f_{y}(y^{\diamond})-f_{y}(y^{\dagger})\rvert+(f_{y}(y^{\dagger})-f_{y}(y^{\circ}))\\ &\leq\left(6\epsilon^{\prime}+4{\epsilon^{\prime}}^{2}+\epsilon^{\prime}(1+\epsilon^{\prime})^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\\ &=\left(7\epsilon^{\prime}+6\epsilon^{\prime 2}+\epsilon^{\prime 3}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (6)
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation ${}^{\prime}$. We have
$f(y^{\diamond})-f_{\inf}$ $\leq(7\epsilon^{\prime}+6\epsilon^{\prime 2}+\epsilon^{\prime 3})\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
$\leq 13\epsilon^{\prime}(1-9\epsilon^{\prime}/2-5{\epsilon^{\prime}}^{2}/4)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
@@
Here, in the first inequality we used (6), the second inequality holds for $\epsilon^{\prime}\in(0,1/13]$, and in the third inequality we used (5). This concludes the first case.
@@
$\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq\frac{2}{\epsilon^{\prime}}p2^{p(p-1)/4}.$

### v2

@@ 5.2 Proof of Proposition 1
In the remainder of the proof, we consider the case $\lVert h\rVert_{2}>\lVert H\rVert_{2}$. In particular, we have $h\neq 0$. Using the algorithm in Lemma 8 (with $\alpha:=1/\lVert h\rVert_{2}^{2}$), we find a positive $l\in\mathbb{Q}$ such that $l\leq 1/\lVert h\rVert_{2}\leq(1+\epsilon)l$. We obtain $l\lVert h\rVert_{2}\leq 1$ and $l\lVert h\rVert_{2}\geq 1/(1+\epsilon)\geq 1-\epsilon$, thus $1-\epsilon\leq l\lVert h\rVert_{2}\leq 1$. Let $\mathcal{B}_{s}:=\mathcal{B}((1-\epsilon)lh,\epsilon)$. From Lemma 9, we either find a vector $s\in\mathcal{B}_{s}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{s})\leq p2^{p(p-1)/4}$. Let $\mathcal{B}_{t}:=\mathcal{B}(-(1-\epsilon)lh,\epsilon)$. From Lemma 9, we either find a vector $t\in\mathcal{B}_{t}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c\})$, or a nonzero vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B}_{t})\leq p2^{p(p-1)/4}$.
Consider the subcase where Lemma 9 successfully found both vectors $s$ and $t$. Note that $\mathcal{B}_{s}$ and $\mathcal{B}_{t}$ are contained in $\mathcal{B}(0,1)$, thus $s,t$ are both feasible to Equation MITR. We apply Lemma 6 and obtain
$f(s)-f(t)\geq 2\left(1-5\epsilon\right)\lVert h\rVert_{2}\geq\left(1-7\epsilon\right)\lVert h\rVert_{2},$
where the last inequality holds for $\epsilon\in(0,1/3]$. Next, consider the subcase where Lemma 9 did not find both vectors $s$ and $t$. In this case, Lemma 9 found a vector $d\in\mathbb{Q}^{n}$ with $d^{\mathsf{T}}b^{1},\dots,d^{\mathsf{T}}b^{p}$ integer and $d^{\mathsf{T}}b^{p+1}=\cdots=d^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d}(\mathcal{B})\leq p2^{p(p-1)/4}$, for $\mathcal{B}\in\{\mathcal{B}_{s},\mathcal{B}_{t}\}$. We then have
$\operatorname{width}_{d}(\mathcal{B}(0,1))$ $=\frac{1}{\epsilon}\operatorname{width}_{d}(\mathcal{B})\leq\frac{1}{\epsilon}p2^{p(p-1)/4}.$
@@
6 Flatness result for Equation E-MIQP
We are now ready to get back to Equation E-MIQP. The main goal of this section is to present a flatness result for this problem: an algorithm that either finds an $\epsilon$-approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. To state this result, and to work with ellipsoids, it will be useful to introduce the following notation. An ellipsoid is a set of the form
@@ 6 Flatness result for Equation E-MIQP
Proposition 2 (Flatness result for Equation E-MIQP).
@@
$\left\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\right\}\subseteq\left[\rho,\rho+19p2^{p(p-1)/4}/\epsilon\right].$
@@
6.1 Ingredients for flatness result
@@
$\begin{split}\min&\quad x^{\mathsf{T}}Hx+h^{\mathsf{T}}x\\ \operatorname{s.t.}&\quad x^{\mathsf{T}}x\leq(1+\delta)^{2}.\end{split}$ (TRδ)
@@
Let $x^{\circ}$ be an optimal solution of Equation TRδ. There is an algorithm which, for any positive $\epsilon\in\mathbb{Q}$, finds a feasible solution $x^{\dagger}$ to Equation TRδ such that
@@
The running time of the algorithm is polynomial in the size of Equation TRδ and in $\log(1/\epsilon)$.
@@
We apply the change of variables $y:=x/(1+\delta)$ to Equation TRδ and obtain
@@
Let $b^{1},\dots,b^{n}$ be linearly independent vectors in $\mathbb{R}^{n}$ and let $p\in\{0,\dots,n\}$. Let $B\in\mathbb{R}^{n\times n}$ with columns $b^{1},\dots,b^{n}$, let $c\in\mathbb{R}^{n}$, and let $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ be the map defined by $\tau(x):=B(x-c)$. Let ${d^{\prime}}\in\mathbb{R}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$. Define $d:=B^{\mathsf{T}}{d^{\prime}}$. Then, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
@@
For $j=1,\dots,n$, we have $d_{j}={b^{j}}^{\mathsf{T}}d^{\prime}$. Thus, ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer implies $d_{1},\cdots,d_{p}$ integer, and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ implies $d_{p+1}=\cdots=d_{n}=0$. We now show that for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, we have $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$
@@ 6.2 Proof of Proposition 2
We now have all ingredients that we need to prove Proposition 2. Before giving the proof, we give an overview of the algorithm. First, we construct an affine transformation $\tau$ that maps $\mathcal{E}(c,Q)$ into an ellipsoid that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,1+\epsilon)$. We denote by Equation E-MIQP the optimization problem obtained by applying the same affine transformation to Equation E-MIQP. Next, we consider Equation E-MIQP-w obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon)$ and by dropping the mixed integer lattice constraint. We then find an almost optimal solution $y^{\dagger}$ to Equation E-MIQP-w and use Lemma 9 to search for a close by vector $y^{\diamond}$ that is feasible to Equation E-MIQP. Next, we consider Equation E-MIQP-s obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP-s that are distant in objective value. If we manage to find all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation E-MIQP. Otherwise, we show that $\mathcal{E}(c,Q)$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 2, which uses Lemmas 10, 11, 9, 1 and 12.
@@
Let $\epsilon^{\prime}:=\epsilon/18\in(0,1/18]$. Using the algorithm in Lemma 10 (with $\delta:=\epsilon^{\prime}$), we find a map $\tau:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\tau(x):=B(x-c)$, with $B\in\mathbb{Q}^{n\times n}$ invertible, such that
@@
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{E}(c,Q))\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},\end{split}$ (E-MIQP)
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider the optimization problem obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the weaker requirement $y\in\mathcal{B}(0,1+\epsilon^{\prime})$ and by dropping the mixed integer lattice constraint:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq(1+\epsilon^{\prime})^{2}.\end{split}$ (E-MIQP-w)
Let $y^{\circ}$ be an optimal solution of Equation E-MIQP-w, which we do not need to compute. Using the algorithm in Lemma 11 (with $\epsilon:=\delta:=\epsilon^{\prime}$), we find a feasible solution $y^{\dagger}$ to Equation E-MIQP-w such that
@@
Let $\mathcal{B}_{\diamond}:=\mathcal{B}((1-\epsilon^{\prime})y^{\dagger}/(1+\epsilon^{\prime}),\epsilon^{\prime})$. From Lemma 9, we either find a vector $y^{\diamond}\in\mathcal{B}_{\diamond}\cap(\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\})$, or a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}_{\diamond})\leq p2^{p(p-1)/4}$.
Consider now the mixed integer trust region problem obtained from Equation E-MIQP by replacing the constraint $y\in\tau(\mathcal{E}(c,Q))$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\}.\end{split}$ (E-MIQP-s)
Using the algorithm in Proposition 1, we either find two feasible solutions $y^{1},y^{2}$ to Equation E-MIQP-s such that
$f(y^{1})-f(y^{2})\geq\left(1-7\epsilon^{\prime}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\},$ (6)
or we find a nonzero vector ${d^{\prime}}\in\mathbb{Q}^{n}$ with ${d^{\prime}}^{\mathsf{T}}b^{1},\dots,{d^{\prime}}^{\mathsf{T}}b^{p}$ integer and ${d^{\prime}}^{\mathsf{T}}b^{p+1}=\cdots={d^{\prime}}^{\mathsf{T}}b^{n}=0$ such that $\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq p2^{p(p-1)/4}/\epsilon^{\prime}$.
In the remainder of the proof, we consider separately two cases. First, consider the case where Lemmas 9 and 1 successfully found all three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$. Note that, since $y^{\dagger}\in\mathcal{B}(0,1+\epsilon^{\prime})$, we have $y^{\dagger}/(1+\epsilon^{\prime})\in\mathcal{B}(0,1)$, thus $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(0,1)$ and $y^{\diamond}$ is feasible to Equation E-MIQP-s. Hence, the three vectors $y^{\diamond}$, $y^{1}$, $y^{2}$ are feasible to Equation E-MIQP-s and to Equation E-MIQP. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation E-MIQP. From (6), we directly obtain the following lower bound on $f_{\sup}-f_{\inf}$:
$\begin{split}f_{\sup}-f_{\inf}&\geq f(y^{1})-f(y^{2})\\ &\geq\left(1-7\epsilon^{\prime}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (7)
Next, we obtain an upper bound on $f_{y}(y^{\diamond})-f_{\inf}$. First, let $g\in\mathbb{R}^{n}$ so that $y^{\diamond}=y^{\dagger}+g$. It is simple to check that $\mathcal{B}_{\diamond}\subseteq\mathcal{B}(y^{\dagger},3\epsilon^{\prime})$, which implies $\lVert g\rVert_{2}\leq 3\epsilon^{\prime}$. Then,
@@
$\leq 6\epsilon^{\prime}\lVert H^{\prime}\rVert_{2}+9{\epsilon^{\prime}}^{2}\lVert H^{\prime}\rVert_{2}+3\epsilon^{\prime}\lVert{h^{\prime}}\rVert_{2}$
$\leq\left(9\epsilon^{\prime}+9{\epsilon^{\prime}}^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.$
@@
$\begin{split}f_{y}(y^{\diamond})-f_{\inf}&\leq f_{y}(y^{\diamond})-f_{y}(y^{\circ})\\ &\leq\lvert f_{y}(y^{\diamond})-f_{y}(y^{\dagger})\rvert+(f_{y}(y^{\dagger})-f_{y}(y^{\circ}))\\ &\leq\left(9\epsilon^{\prime}+9{\epsilon^{\prime}}^{2}+\epsilon^{\prime}(1+\epsilon^{\prime})^{2}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\\ &=\left(10\epsilon^{\prime}+11\epsilon^{\prime 2}+\epsilon^{\prime 3}\right)\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}.\end{split}$ (8)
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation E-MIQP. We have
$f(y^{\diamond})-f_{\inf}$ $\leq(10\epsilon^{\prime}+11\epsilon^{\prime 2}+\epsilon^{\prime 3})\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
$\leq 18\epsilon^{\prime}(1-7\epsilon^{\prime})\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}$
@@
Here, in the first inequality we used (8), the second inequality holds for $\epsilon^{\prime}\in(0,1/18]$, and in the third inequality we used (7). It then follows that $x^{\diamond}:=B^{-1}y^{\diamond}+c$ is an $\epsilon$-approximate solution to Equation E-MIQP. This concludes the first case.
@@
$\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1))\leq\frac{1}{\epsilon^{\prime}}p2^{p(p-1)/4}.$